Find the discriminant then tell whether the equation has two,one, or no solutions
1.3x^2+4x-2=0
2.-5x^2+6x-6=0
3.6x^2-2x+4=0

Answer by Maths68(1474)

(1)
3x^2+4x-2=0
a=3, b=4, c=-2
Discriminant = b^2-4ac
Discriminant = (4)^2-4(3)(-2)
Discriminant = 16+24
Discriminant = 40
Discriminant > 0 Therefore there are two solutions.


(2)
-5x^2+6x-6=0
a=-5, b=6, c=-6
Discriminant = b^2-4ac
Discriminant = (6)^2-4(-5)(-6)
Discriminant = 36-120
Discriminant = -84
Discriminant < 0 Therefore there is no real number solution

(3)
6x^2-2x+4=0
2(3x^2-x+2)=0
2(3x^2-x+2)/2=0/2
3x^2-x+2=0
Discriminant = b^2-4ac
a=3, b=-1, c=2
Discriminant = (-1)^2-4(3)(2)
Discriminant = 1-24
Discriminant = -23
Discriminant < 0 Therefore there is no real number solution